A pet store has 7 puppies, including 4 poodles, 2 terriers, and 1 retriever. If Rebecca and Aaron in that order each select one puppy at random, with replacement, what is the probability they a) both select a poodle, b) Rebecca selects a retriever, Aaron selects a terrier, c) Rebecca selects a terrier, Aaron selects a retriever, d) both select a retriever?
Answer by mananth(16949)


a pet store has seven puppies including 4 poodles, 2 terriers, and 1 retriever. If Rebecca and Aaron, in that order, each select one puppy at random, with replacement (they may both select the same one),

find the probability that
a) both select a poodle===> (4/7)^2

b) Rebecca selects a retriever and Aaron selects a terrier
====>(1/7)(2/7) = 2/49


c) Rebecca selects a terrier and Aaron selects a retriever
===> (2/7)(1/7) = 2/49
d) both select a retriever::: (1/7)^2 = 1/49